OpenStudy (rmrjr22):

Given the curve y=(9-3x)^1/2 Find the volume formed when the curve is rotated about the x-axis from x=0 to x=3....

OpenStudy (campbell_st):

well \[V = \pi \int\limits_{a}^{b} y^2 dx\] so you have \[V = \pi \int\limits_{0}^{3} ((9 - 3x)^{1/2})^2 dx\]

OpenStudy (campbell_st):

\[so you have V = \pi \int\limits_{0}^{3} 9 - 3x dx\]

OpenStudy (rmrjr22):

Would this be the Shell Method?

OpenStudy (campbell_st):

well is very straight forward \[V = \pi [ 9x - 3/2x^2] \] just evaluate it now f(3) - f(0) might be 13.5 pi units cubed

OpenStudy (rmrjr22):

I ended up getting 42.412 units

OpenStudy (campbell_st):

well 13.5 x 3.14 is about the 43 as I don't have a calculator with me

OpenStudy (rmrjr22):

There was another part to it... Compare this to the volume of the circular cylinder formed by rotating the line y=3 over the same interval about the x-axis. What is the relationship between the 2 volumes?

OpenStudy (campbell_st):

well the cylinder is easy... the radius is 3... distance from y = 0 (x axis) to y = 3) height of the cylinder is 3 units... distance between x = 0 and x = 3 use \[V = \pi r^2 h\]

OpenStudy (campbell_st):

I think the cylinder is larger

OpenStudy (campbell_st):

actually the cylinder is double the size.. \[V _{cylinder} = 27 \pi \] \[V _{curve} = 13.5 \pi\]

OpenStudy (rmrjr22):

\[(3.14)\int\limits_{0}^{3}(3x)^2dx \] ?

OpenStudy (rmrjr22):

No... Thats not it...